# ROLCH: Regularized Online Learning for Conditional Heteroskedasticity

### v1

@@ 2.3 Online GAMLSS
and can only update $w_{kk,n+1}^{[m,r]}$. Note that due to the update rules in Equations 26 and 27, this effect is counter-weighted by the exponential forget $\gamma$. However, there is a delicate balance to strike to trade off the beneficial effect of the forget and increased instability in the coefficient estimation. This might lead to slower convergence compared to batch learning if the data is drawn from a stationary process. The online GAMLSS is, therefore, an approximation of the repeated batch GAMLSS, contrary to the case for, e.g. recursive least squares or online coordinate descent, where the update step leads to the exact coefficients.
@@ 3 Implementation in the Python Package ROLCH
  • 1.
    (Consistent API) Our package provides an API comparable to other major machine learning packages in Python such as sklearn or tensorflow with very few dependencies (only numpy, numba, and scipy Harris et al., 2020; Virtanen et al., 2020; Lam et al., 2015).
  • 2.
    (Automatic scaling) We employ automatic mean-variance scaling for all covariates with an incremental calculation of the exponentially discounted mean and variance using Welford’s Algorithm (Welford, 1962).
  • 3.
    (Online Model selection) We extend the model selection approach in Ziel et al. (2021) and allow for online automatic model selection using AIC, BIC, and HQC allowing for fast, efficient and theoretically grounded (local) model selection in the online coordinate descent.
  • 4.
    (Performance) We employ just-in-time compilation using numba to achieve a high-performance implementation and employ various computational tricks such as active set coordinate descent, different warm-starting options (using the previous fit or a mixture of previous fit and previous $\beta_{\lambda}$ on the same coefficient path) and allow for random selection during coordinate descent to speed up the convergence of the coordinate descent (see e.g. Shi et al., 2016; Wright, 2015).
  • 5.
    (Flexibility) At the same time, our implementation offers multiple features for users such as choosing the set of regularized coefficients and the possibility to constrain coefficients within bounds, e.g. allow for positive/negative coefficients only. The update formula for the coordinate descent (Equation 22) is then
@@
Currently, our package implements the Gaussian distribution, Student’s $t$-distribution, and Johnson’s $S_{U}$ distribution. [^4] Since the implementations derive from scipy, implementing further distributions is straightforward. Only the cross-derivatives and the initial guess for the $\theta_{k}^{[0,0]}$ need to be added manually. The code for our implementation can be found at https://github.com/simon-hirsch/rolch and the package is available at the Python Package Index (https://pypi.org/project/rolch/) under the MIT-License.
4 Forecasting Study for Electricity Price Forecasting
The following section presents an exemplary application in electricity price forecasting (EPF) for the online GAMLSS method. We employ the same setting as Marcjasz et al. (2023) in forecasting day-ahead electricity prices for the German short-term electricity market.
@@ 4 Forecasting Study for Electricity Price Forecasting
The German electricity market consists of four major markets: The long-term futures market, the day-ahead spot market, the intraday market and the balancing market. Since our forecasting study is concerned with the day-ahead or spot market only, our description focuses on this market only. A detailed summary of the German electricity market can be found in Viehmann (2017). The spot market is the major reference price for the futures market and its trading volume is a multiple of the intraday and balancing market. Let $h\in\{0,1,...,H\}$ and $H=23$ denote the 24 delivery hours and $d$ denote the delivery day. The market is organized as a daily auction at $d-1$, 12:00 hours for all 24 delivery hours of the following day. Results are published at $d-1$, 12:42 hours. Market participants submit limit orders and block/linked/complex bids. The market is organized as a pay-as-cleared auction and cleared by the EUPHEMIA algorithm and maximises the European Welfare under consideration of constraints such as the available international transmission capacity (Viehmann, 2017; Marcjasz et al., 2023; Uniejewski & Weron, 2021; Ziel & Weron, 2018; Lago et al., 2018, 2021).
@@
We employ a (linear) expert model type as it is common in electricity price forecasting Lago et al. (2018); Marcjasz et al. (2023); Ziel & Weron (2018). The model consists of 38 terms and captures the autoregressive price effects, seasonal effects, the fundamental effects of renewable generation and the influence of fuel prices. In the distributional framework, we model each distribution parameter $k=1,...,p$ for each delivery hour $h=\{0,1,...,H\}$ individually as function of linear predictors:
@@
We benchmark forecasts using established, proper probabilistic scoring rules (Gneiting, 2008, 2011; Gneiting & Katzfuss, 2014; Nowotarski & Weron, 2018). For the mean prediction, we employ the Root Mean Squared Error (RMSE) and the Mean Absolute Error (MAE). We evaluate the Coverage (CR) and the Interval Score (IS, also known as Winkler Score, see Bracher et al. (2021)) for the 50%, 75%, 90% and 95% prediction intervals (PI). For the full predictive distribution, we evaluate the Log Score (LS) and the continuous ranked probability score (CRPS) using the approximation via the Pinball Score (PS) on a dense grid of quantiles ${\mathcal{Q}}=\{0.01,0.02,0.03,...,0.99\}$. The implementation of the scoring rules is provided by the scoringrules package (Zanetta & Allen, 2024). We evaluate the statistical significance of the difference in predictive accuracy using the Diebold-Mariano test Diebold & Mariano (2002); Diebold (2015).
@@
Furthermore, we report the Log Score (LS) as the negative log-likelihood of the true value under the predictive distribution. The LS is defined as:
@@
where $\widehat{\boldsymbol{d}}_{P_{d,h}}(x)$ is the probability density function of the predictive distribution $\widehat{{\mathcal{D}}}^{P}_{d,h}$ and observation $x$. Let us note that the RMSE, MAE, IS, LS and the CRPS [^5] Furthermore, note that within the EPF community, the first fraction in the formulation of the CRPS is often $1/|{\mathcal{Q}}|$ instead of $2/|{\mathcal{Q}}|$. This formulation corresponds to $0.5\times\operatorname{CRPS}$ and has initially been used this way in the GEFcom 2014 as the Average Pinball Score (Hong et al., 2016; Marcjasz et al., 2023; Nowotarski & Weron, 2018). are strictly proper scoring rules. For all of them, lower scores correspond to a better forecast. The CR is a measure of calibration for probabilistic forecasts only. Lastly, we evaluate the computation time for all forecasting studies.
@@
5 Results and Discussion
Table 1 presents the tested models in the different settings, estimation methods and the results of the different scoring rules. Figure 1 gives the $p$-value of the pairwise Diebold-Mariano test. In our case study, online learning models deliver competitive performance and, to a certain extent, can even outperform rolling batch estimation while reducing the estimation time by some 1-2 orders of magnitude. The following paragraphs discuss some observations in more detail.
MAE RMSE CR50 CR90 IS50 IS90 LS CRPS Time (Min)
@@ 5 Results and Discussion
Method Distribution Estimation Location Scale Tail Skew
Batch Student-$t$ OLS Eq. (29) Eq. (29) - - 4.400 7.473 0.708 0.961 15.450 27.022 3.259 3.451 26.296
LASSO Eq. (29) Eq. (29) - - 5.078 8.834 0.510 0.881 16.961 22.787 3.469 3.850 167.578
Johnson-$S_{U}$ OLS Eq. (29) - - - 4.529 7.139 0.555 0.916 14.910 21.434 3.340 3.371 458.131
Eq. (29) Eq. (29) Eq. (29) 4.158 9.163 0.465 0.850 13.558 17.255 3.163 3.057 1670.944
Gaussian $\mathcal{N}$ OLS Eq. (29) - - - 4.649 6.956 0.784 0.982 17.156 33.637 3.476 3.780 13.177
Eq. (29) - - 4.634 6.949 0.689 0.963 16.087 28.445 3.349 3.578 27.702
LASSO Eq. (29) Eq. (29) - - 9.596 14.243 0.589 0.912 31.841 47.327 4.032 7.154 479.687
Online Student-$t$ OLS Eq. (29) - - - 4.262 6.980 0.557 0.920 14.164 21.146 3.165 3.217 1.610
Eq. (29) - - 4.216 7.034 0.533 0.905 13.957 20.041 3.122 3.153 1.910
Eq. (29) - 4.239 7.073 0.591 0.914 15.058 24.145 3.256 3.348 1.904
LASSO Eq. (29) - - - 4.257 6.980 0.557 0.921 14.151 21.126 3.164 3.214 2.790
Eq. (29) - - 4.285 7.027 0.550 0.932 14.598 24.986 3.163 3.304 3.266
Eq. (29) - 4.296 7.063 0.514 0.914 14.527 21.730 3.183 3.274 10.291
Johnson-$S_{U}$ OLS Eq. (29) - - - 4.396 7.147 0.561 0.912 14.566 20.888 3.326 3.294 2.170
Eq. (29) - - 4.328 7.556 0.510 0.877 14.290 19.625 3.183 3.227 3.051
Eq. (29) - 4.294 9.342 0.455 0.836 14.075 17.224 3.190 3.183 3.742
Eq. (29) 4.388 9.360 0.470 0.850 14.416 18.528 3.160 3.259 4.270
Gaussian $\mathcal{N}$ OLS Eq. (29) - - - 4.486 6.871 0.654 0.951 15.198 24.958 3.384 3.414 1.015
Eq. (29) - - 4.282 6.954 0.625 0.926 19.336 34.899 3.390 4.086 1.237
@@
Table 1: Models and Scores. Combinations of models analysed in the forecasting study. Distribution refers to the assumed parametric form of the response distribution. Setting refers to whether models are estimated incrementally or re-trained on the increasing window training set. The best value in each column is marked bold. Note that the timing corresponds to a full forecasting study, i.e. the estimation of $736\times 24=17.664$ models. The colour gradient runs on a logarithmic scale.
@@
Figure 2: Illustrative Forecast. Prediction Intervals derived from the probabilistic model. Here, we assume the power prices to follow Johnson’s $S_{U}$ distribution and model all moments conditionally on Equation 29. The prediction intervals correspond to the $\{0.05,0.1,...,0.95\}$ quantiles of the predictive distribution. Note that the period with extremely low prices corresponds to Christmas.
Normal Distribution
The normal distribution is common as a simple benchmark and serves this purpose also in this forecasting study. However, in the GAMLSS framework (in the batch and online setting, in the R and our Python implementation), we have encountered issues with the estimation of models where the specified distribution does not match the empirical distribution of the data (see e.g. Robledo, 2022). Since electricity prices are clearly non-Gaussian, this also applies in this case.
LASSO estimation
We note that LASSO estimation does not necessarily improve the performance of the model. This can be seen, e.g. in the direct comparison between the location-scale $t$ model in the online and batch settings. This is in contrast to the literature, where LASSO estimated models typically outperform OLS models (Uniejewski & Weron, 2021; Nowotarski & Weron, 2018). However, these results hold primarily for models in the expected value and little research is done on model selection in the higher moments. Additionally, we have not (yet) performed any tuning with respect to the model selection, but estimated all models taking the BIC.
6 Conclusion
Methodological contribution
This paper presents an efficient, scalable approach to online distributional regression respectively online regression for conditional heteroskedastic data. We implement an incremental algorithm for the well-known, linear-parametric GAMLSS modelling approach, including regularized estimation Rigby & Stasinopoulos (2005); Stasinopoulos & Rigby (2008); Stasinopoulos et al. (2018); Klein (2024) and provide an open-source Python implementation of our approach.

### v2

@@ 3 Online Distributional Regression
and can only update $w_{k,k,n+1}^{[m,r]}$. Note that due to the update rules in Equations 22 and 23, this effect is counter-weighted by the exponential forget $\gamma$. However, there is a delicate balance to strike to trade off the beneficial effect of the forget and increased instability in the coefficient estimation. This might lead to slower convergence compared to batch learning if the data is drawn from a stationary process. The online GAMLSS is, therefore, an approximation of the repeated batch GAMLSS, contrary to the case for, e.g. recursive least squares or online coordinate descent, where the update step leads to the equivalent results.
When choosing a specific distribution assumption for ${\mathcal{D}}$ in the GAMLSS framework, the weight matrix for $\theta_{k}$ in each update step can recovered with the help of Equations 3 - 6. The online weight matrix is retrieved by iteratively inserting the weight matrix in Equation (25). For illustration, building on the standard linear-Gaussian case for the GAMLSS in Proposition 1 we can examine the difference explicitly for the mean equation:
Corollary 1
Assume the linear-Gaussian case for the GAMLSS with $\theta_{1}=\mu$ and $\theta_{2}=\sigma^{2}$ of Proposition 1. In the batch and online setting with initial batch $n_{0}<n$, the weight matrices for $\theta_{1}=\mu$ for $n+1$ are:
${\mathcal{W}}_{1,1,n+1}^{[m,r],\operatorname{online}}$ $=\operatorname{diag}\Biggl{(}\underbrace{\dfrac{\gamma^{n}}{\hat{\sigma}^{2}_{0|n_{0}}},...,\dfrac{\gamma^{n-n_{0}}}{\hat{\sigma}^{2}_{n_{0}|n_{0}}}}_{\text{Estimates initial batch $n_{0}$}},\dfrac{\gamma^{n-n_{0}-1}}{\hat{\sigma}^{2}_{n_{0}+1|n_{0}+1}},...,\dfrac{\gamma}{\hat{\sigma}^{2}_{n|n}},\dfrac{1}{(\hat{\sigma}^{2}_{n+1|n+1})^{[m,r]}}\Biggl{)}$
${\mathcal{W}}_{1,1,n+1}^{[m,r],\operatorname{batch}}$ $=\operatorname{diag}\left(\dfrac{\gamma^{n}}{(\hat{\sigma}^{2}_{0|n+1})^{[m,r]}},...,\dfrac{\gamma^{n-n_{0}-1}}{(\hat{\sigma}^{2}_{n_{0}|n+1})^{[m,r]}},...,\dfrac{\gamma}{(\hat{\sigma}^{2}_{n|n+1})^{[m,r]}},\dfrac{1}{(\hat{\sigma}^{2}_{n+1|n+1})^{[m,r]}}\right)$
where $\gamma=(1-\varphi)$ is the exponential discounting with forgetting factor $\varphi$. The difference can be summarized in
${\mathcal{W}}_{1,1,n+1}^{[m,r],\operatorname{online}}-{\mathcal{W}}_{1,1,n+1}^{[m,r],\operatorname{batch}}=\sum_{i=0}^{n_{0}}\dfrac{\gamma^{n-i}}{\hat{\sigma}^{2}_{i|n_{0}}}-\dfrac{\gamma^{n-i}}{(\hat{\sigma}^{2}_{i|n+1})^{[m,r]}}+\sum_{i=n_{0}+1}^{n}\dfrac{\gamma^{n-i}}{\hat{\sigma}^{2}_{i|i}}-\dfrac{\gamma^{n-i}}{(\hat{\sigma}^{2}_{i|n+1})^{[m,r]}}.$ (26)
From Corollary 1 we see that the difference in the online and batch setting depends on the length of the initial batch $n_{0}$ and the strength of the exponential discounting. This result differs from the recursive least squares setting, where the batch and online results coincide. In the setting of Corollary 1, we can similarly state the weight matrix of ${\mathcal{W}}_{2,2}$ for $\theta_{2}=\sigma^{2}$. More complex weight matrices can be derived by utilizing different combinations of the link function $g(\cdot)$ and distribution ${\mathcal{D}}$.
Mini-batch update steps
The online GAMLSS Algorithm 4 implements one-step updates, i.e. receiving new data row-by-row: in each update, we receive $\boldsymbol{X}_{k,n+1}$ and $\boldsymbol{Y}_{k,n+1}$. In principle, mini-batch update steps, in which we receive ${\mathcal{X}}_{k,n+1:n+t}=(\boldsymbol{X}_{k,n+1},...,\boldsymbol{X}_{k,n+t})^{\top}$ and ${\mathcal{Y}}_{n+1:n+t}=(Y_{n+1},...,Y_{n+t})$ in a single pass can be implemented in the same manner. Equations (22) and (23) natively handle mini-batch updates. Equation (24) needs to be adjusted to
$\text{RSS}_{k,n+t,\lambda}=\dfrac{\left((1-\varphi)^{t-1}\boldsymbol{W}_{k,k,n+t}\left(\boldsymbol{z}_{n+t,k}^{[m,r]}-\boldsymbol{\beta}_{n+t,k,\lambda}^{[m,r]}{\mathcal{X}}^{\top}_{k,n+1:n+t}\right)\right)^{2}+(1-\varphi)^{t}\text{RSS}_{n,k,\lambda}}{\sum_{i=0}^{t-1}(1-\varphi)^{i}w_{k,k,n+i}+(1-\varphi)^{t}\omega_{k,k,n}}$
@@ 4 Implementation in the Python Package ROLCH
  • Our package provides an API comparable to other major machine learning packages in Python such as sklearn or tensorflow with very few dependencies (only numpy, numba, and scipy, see Harris et al., 2020; Virtanen et al., 2020; Lam et al., 2015).
  • We employ automatic mean-variance scaling for all covariates with an incremental calculation of the exponentially discounted mean and variance using Welford’s Algorithm (Welford, 1962).
  • We extend the model selection approach in Ziel et al. (2021) and allow for online automatic model selection using AIC, BIC, and HQC allowing for fast, efficient and theoretically grounded (local) model selection in the online coordinate descent.
  • We employ just-in-time compilation using numba to achieve a high-performance implementation and employ various computational tricks such as active set coordinate descent, different warm-starting options (using the previous fit or a mixture of previous fit and previous $\beta_{\lambda}$ on the same coefficient path) and allow for random selection during coordinate descent to speed up the convergence of the coordinate descent (see e.g. Shi et al., 2016; Wright, 2015).
  • Our implementation offers multiple features for users such as choosing the set of regularized coefficients and the possibility to constrain coefficients within bounds, e.g. allow for positive/negative coefficients only. The update formula for the coordinate descent (equation (18)) is then
@@
Currently, our package implements the Gaussian distribution, Student’s $t$-distribution, and Johnson’s $S_{U}$ distribution. [^4] Since the implementations derive from scipy, implementing further distributions is straightforward. Only the cross-derivatives and the initial guess for the $\theta_{k}^{[0,0]}$ need to be added manually. The code for our implementation is open source (see https://github.com/simon-hirsch/rolch) and the package is available at the Python Package Index (see https://pypi.org/project/rolch/) under the MIT-License.
5 Simulation Study
The following section presents a small simulation study on the approximation quality of the online GAMLSS model to analyse the impact of the length of the initial training set and the forget. Our main metric is the $\ell_{1}$-norm of the difference between the online updated coefficients and the true coefficients in comparison to regular refitted batch models.
Setting
We consider a setting where ${\mathcal{Y}}_{i}$ is normally distributed where the location and scale parameter depend on the entries of the covariate matrices ${\mathcal{X}}_{\mu}$ and ${\mathcal{X}}_{\sigma}$ of dimension $J$. We simulate the entries of ${\mathcal{X}}$ by a $J$-dimensional standard normal distribution ${\boldsymbol{X}_{i}\sim{\mathcal{N}}(0_{J},I_{J})}$ and define
$Y_{i}\sim{\mathcal{N}}\left(\boldsymbol{X}_{i,\mu}\beta_{\mu},\;\operatorname{exp}(\boldsymbol{X}_{i,\sigma}\beta_{\sigma})\right).$ (27)
The $J$-dimensional coefficient vectors $\beta_{0}$ and $\beta_{1}$ are drawn from the uniform distribution except that $K$ positions are fixed to zero:
$\beta_{\mu}\sim\delta_{0}^{\otimes{K}}\otimes{\mathcal{U}}(-1,1)^{\otimes{J-K}}\text{ and }\beta_{\sigma}\sim\delta_{0}^{\otimes{K}}\otimes{\mathcal{U}}(0,\log(2)/4)^{\otimes{J-K}},$ (28)
where the exponential ensures that the standard deviation is positive. Accordingly, we use a log-link for the scale parameter of the normal distribution. We use $J=21$ and set five of them as uninformative features ($K=5$). We allow for an initial batch of $n_{0}=\{400,1600,6400\}$ and update the online GAMLSS model each step. On the quadratic grid $G=\{400,900,1600,...,25600\}$, we compare the online update coefficients and a batch fit with the true coefficients by evaluating the $\ell_{1}$-norm $\|\cdot\|_{1}$ between the true and estimated coefficients. Corresponding with the results of Corollary 1 that higher forgets should decrease the difference between batch and online estimation, we additionally estimate three models with forget $\varphi=\{0.001,0.0001,0.00001\}$ starting from an initial batch $n_{0}=400$.
Figure 1: Simulation study results: The difference $\|\beta^{\text{True}}_{k}-\hat{\beta}_{k}\|_{1}$ relative to the selected initial batch size $n_{0}$. The main panel shows box-plots all $M=100$ simulations. The inner panel shows the difference in the means for the repeated batch and online estimation for the same initial batch size $n_{0}$ at the selected grid points. The top plot gives results for $\theta_{0}=\mu$, the bottom plot gives results for $\theta_{1}=\sigma$.
Figure 2: Simulation study results: The difference $\|\beta^{\text{True}}_{k}-\hat{\beta}_{k}\|_{1}$ relative to the selected forget factor $\varphi$. The main panel shows box-plots all $M=100$ simulations. The inner panel shows the difference in the means for the repeated batch and online estimation for the same forget $\varphi$ at the selected grid points. The top plot gives results for $\theta_{0}=\mu$, the bottom plot gives results for $\theta_{1}=\sigma$.
Results
Figures 1 and 2 give the results of the simulation study with respect to the initial batch size $n_{0}$ and the forget $\varphi$. Generally, the results are aligned with Corollary 1. We see that the difference between the online setting and the repeated batch estimation decreases as more observations are consumed. We see lower errors for larger initial batch sizes also a faster decrease of the error. Additionally, we see in Figure 2 that models having a small forget generally converge faster towards the batch model. Note that the effective training size of the model taking the smallest forget is $1000001/\varphi_{0}=$ 100000, but the difference between repeated batch estimation and online estimation vanishes after 25.600 observations already. We also see that the model with the highest forget $\varphi=0.01$ and an effective training size of 100 observations retains a consistent error in the estimation of the scale parameter $\theta_{1}=\sigma$.
6 Forecasting Study for Electricity Prices
The following section presents an exemplary application in electricity price forecasting (EPF) for the online GAMLSS method. We employ the same setting as Marcjasz et al. (2023) in forecasting day-ahead electricity prices for the German short-term electricity market. An illustrative forecast can be seen in Figure 3.
@@ 6 Forecasting Study for Electricity Prices
The German electricity market consists of four major markets: The long-term futures market, the day-ahead spot market, the intraday market and the balancing market. Since our forecasting study is concerned with the day-ahead or spot market only, our description focuses on this market only. The spot market is the major reference price for the futures market and its trading volume is a multiple of the intraday and balancing market. Let $h\in\{0,1,...,H\}$ and $H=23$ denote the 24 delivery hours and $d$ denote the delivery day. The market is organized as a daily auction at $d-1$, 12:00 hours for all 24 delivery hours of the following day (further information can be found in, e.g. Viehmann, 2017; Marcjasz et al., 2023; Uniejewski and Weron, 2021; Ziel and Weron, 2018; Lago et al., 2018, 2021).
@@
We employ a (linear) expert model type as it is common in electricity price forecasting Lago et al. (2018); Marcjasz et al. (2023); Ziel and Weron (2018). The model consists of 38 terms and captures the autoregressive price effects, seasonal effects, the fundamental effects of renewable generation and the influence of fuel prices. We employ the Student-$t$ and Johnson’s $S_{U}$ (JSU) distribution as both have been used for electricity price forecasting and other financial applications. In the distributional framework, we model each distribution parameter $k=1,...,p$ for each delivery hour $h=\{0,1,...,H\}$ individually as function of linear predictors:
@@
Figure 3: Illustrative Forecast. Prediction Intervals derived from the online probabilistic model. Here, we assume the power prices to follow Johnson’s $S_{U}$ distribution and model all moments conditionally on Equation 29. The prediction intervals correspond to the $\{0.01,0.05,0.1,...,0.95,0.99\}$ quantiles of the predictive distribution. Note that the period with extremely low prices corresponds to Christmas.
@@
We benchmark forecasts using established, proper probabilistic scoring rules (Gneiting, 2008, 2011; Gneiting and Katzfuss, 2014; Nowotarski and Weron, 2018). For the mean prediction, we employ the Root Mean Squared Error (RMSE) and the Mean Absolute Error (MAE). We evaluate the Coverage (CR) and the Interval Score (IS, also known as Winkler Score, see Bracher et al. (2021)) for the 50%, 75%, 90% and 95% prediction intervals (PI). For the full predictive distribution, we evaluate the Log Score (LS) and the continuous ranked probability score (CRPS) using the approximation via the Pinball Score (PS) on a dense grid of quantiles ${\mathcal{Q}}=\{0.01,0.02,0.03,...,0.99\}$. The implementation of the scoring rules is provided by the scoringrules package (Zanetta and Allen, 2024). We evaluate the statistical significance of the difference in predictive accuracy using the Diebold-Mariano test Diebold and Mariano (2002); Diebold (2015).
@@
Furthermore, we report the Log-Score (LS) as the negative log-likelihood of the true value under the predictive distribution. The LS is defined as:
@@
where $\widehat{\boldsymbol{d}}_{P_{d,h}}(x)$ is the probability density function of the predictive distribution $\widehat{{\mathcal{D}}}^{P}_{d,h}$ and observation $x$. Let us note that the RMSE, MAE, IS, LS and the CRPS [^5] Furthermore, note that within the forecasting community, the CRPS is sometimes reported using $1/|{\mathcal{Q}}|$ as the first fraction instead of $2/|{\mathcal{Q}}|$. This formulation corresponds to $0.5\times\operatorname{CRPS}$ and has initially been used this way in the GEFcom 2014 as the Average Pinball Score (APS) (Hong et al., 2016; Marcjasz et al., 2023; Nowotarski and Weron, 2018). are strictly proper scoring rules. For all of them, lower scores correspond to a better forecast. The CR is a measure of calibration for probabilistic forecasts only. Lastly, we evaluate the computation time for all forecasting studies.
@@
    Method     Distribution     Estimation     Location     Scale     Tail     Skew     MAE     RMSE     CR50     CR90     IS50     IS90     LS     CRPS     Time (Min)
Batch t OLS - - 4.30 7.26 0.46 0.87 14.26 19.08 3.09 3.22 122.31
Batch t LASSO - - 4.26 7.23 0.46 0.87 14.12 18.83 3.07 3.19 306.27
Batch JSU OLS - - - 4.53 7.13 0.56 0.92 14.92 21.51 3.34 3.37 332.42
Batch JSU OLS 4.32 8.73 0.48 0.87 14.01 18.21 3.18 3.16 1302.20
Batch JSU LASSO - - - 4.48 7.11 0.57 0.92 14.77 21.53 3.33 3.34 456.18
Batch JSU LASSO 4.17 8.88 0.48 0.87 13.54 17.71 3.15 3.05 1787.15
Online t OLS - - - 4.35 6.98 0.55 0.94 14.61 25.74 3.21 3.33 1.00
Online t OLS - - 4.30 7.32 0.47 0.88 14.29 19.36 3.08 3.23 1.17
Online t OLS - 4.32 7.29 0.51 0.88 15.07 23.18 3.80 3.36 1.47
Online t LASSO - - - 4.35 6.98 0.55 0.93 14.65 25.76 3.21 3.34 1.06
Online t LASSO - - 4.29 7.31 0.47 0.88 14.23 19.36 3.07 3.21 1.92
Online t LASSO - 4.29 7.30 0.48 0.88 14.20 19.74 3.10 3.20 3.82
Online JSU OLS - - - 4.40 7.17 0.56 0.92 14.57 21.00 3.33 3.30 1.34
Online JSU OLS - - 4.33 7.53 0.51 0.88 14.29 19.60 3.18 3.23 1.89
Online JSU OLS - 4.29 9.31 0.46 0.84 14.08 17.25 3.19 3.18 2.29
Online JSU OLS - 4.36 7.44 0.51 0.88 14.44 20.06 3.21 3.26 2.23
Online JSU OLS 4.37 8.83 0.49 0.87 14.38 19.19 3.18 3.25 2.46
Online JSU LASSO - - - 4.37 7.11 0.57 0.92 14.50 21.03 3.32 3.28 1.67
Online JSU LASSO - - 4.30 7.55 0.51 0.88 14.19 19.45 3.16 3.20 2.31
Online JSU LASSO - 4.24 9.15 0.46 0.84 13.88 17.06 3.17 3.14 4.35
Online JSU LASSO - 4.33 7.48 0.52 0.89 14.35 20.47 3.19 3.24 3.15
Online JSU LASSO 4.43 8.80 0.49 0.87 14.48 18.91 3.20 3.27 4.92
@@
Table 1: Models and Scores. Combinations of models analysed in the forecasting study. Distribution refers to the assumed parametric form of the response distribution. Setting refers to whether models are estimated incrementally or re-trained on the increasing window training set. The best value in each column is marked bold. Note that the timing corresponds to a full forecasting study, i.e. the estimation of $736\times 24=17.664$ models.
Results and Discussion
Table 1 presents the tested models in the different settings, estimation methods and the results of the different scoring rules. Figure 4 gives the $p$-value of the pairwise Diebold-Mariano test. In our case study, online learning models deliver competitive performance and, to a certain extent, can even outperform rolling batch estimation while reducing the estimation time by 2-3 orders of magnitude. The following paragraphs discuss some observations in more detail.
  • On a high level, the difference in predictive performance in terms of the CRPS between the pairs of repeated batch estimation and online models is often marginal and statistically not significant (see Figure 4, e.g. the comparison between batch and online JSU models with 3 and 4 parameters modelled and the location-scale student-$t$ model). Again, we attribute the differences in the performance primarily to the results of Proposition 1 and Corollary 1, which can be translated to the Student-$t$ and JSU case.
  • In terms of the CRPS, the best-performing online model is the LASSO-JSU model, including equations for the conditional location, scale and skewness parameters. For the repeated batch estimation, the best performance is achieved by the LASSO-Student-$t$ model. Models such as the 4-parameter online OLS-JSU deliver close performance.
  • Let us note that most of the employed covariates have high, fundamentally motivated explanatory power (Nowotarski and Weron, 2018; Marcjasz et al., 2023; Ziel and Weron, 2018), which explains the small performance gap between the LASSO and OLS estimated models. Nevertheless, we note that regularisation improves the predictive accuracy, especially for models where the tail behaviour and skewness are modelled, which aligns with general results on distributional modelling (see e.g. Ziel (2022) and the paragraph on error propagation in Section 2.1).
  • In terms of computation time (see Table 1, last column), the online models are estimated 2-3 orders of magnitude faster. However, comparing the estimation time between OLS and LASSO for batch and online models, we note that we experience a higher relative increase in estimation time when using LASSO in the online setting. Conversely, the relative increase in estimation time for modelling more distribution parameters decreases in the online setting.
  • Lastly, our results are competitive with other results on the same dataset. Marcjasz et al. (2023) report average pinball scores (APS) of 1.575 to 1.662 (equiv. CRPS of 3.150 to 3.324) for batch linear distributional models using quantile regression averaging and of Brusaferri et al. (2024), who report APS of 1.450 to 1.549 (CRPS of 2.900 to 3.098) for adaptive conformal prediction methods, however, the underlying model is based on batch, non-linear neural networks which provide lower MAEs as well, translating towards lower distributional scores as well.
@@
Wraping up, our results show that the online GAMLSS provides competitive forecasting performance, both towards the repeated batch estimation and other probabilistic prediction methods at significantly reduced computation time.
7 Discussion and Conclusion
Contribution
This paper presents an efficient, scalable approach to online distributional regression. We implement an incremental approximation algorithm for the well-known, linear-parametric GAMLSS (for batch settings see, e.g. Rigby and Stasinopoulos, 2005; Klein, 2024). We employ regularized estimation and provide an approach for online model selection. Furthermore, we explicitly show the relationship between IRLS for conditional heteroskedasticity (Dette and Wagener, 2013; Ziel, 2016) in the batch setting and use it to exemplary show the limitations of the online GAMLSS. Lastly, we provide an open-source Python implementation of our approach.